A plane travels at 700 mph. How long does it take for the plane to fly 4 miles?
Answer by mananth(16949)

speed = 700 mph
distance= 4 miles

Time = distance/speed
Time = 4/700=0.00571 hours
=0.00571 * 3600 = 20.6 seconds